autumnspivey:

Jordan solved the equation −7x + 25 = 48; his work is shown below. Identify the error and where it was made.

lilaspie7:

do u have a screenshot of him solving the problem?

XioGonz:

He/She wrote \[\frac{ 7*X }{ 7 }=\frac{ 23 }{ 7 }\] He/She was supposed to write \[\frac{ -7*X }{ -7 }=\frac{ 23 }{ -7 }\] So his/her answer was \[\frac{ 23 }{ 7 }\]It was supposed to be \[-\frac{ 23 }{ 7 }\]

CripQUEZZ:

Ok so first u gotta get rid of 25 so -7x + 25 = 48 - 25 - 25= 23 Then do the opposite of multiplying 7x 23 — 7 And since ur gonna get a decimal it stays 23 — -7